Pamela is 10 years younger than Jiri. The sum of their ages is 64. What is Jiris age?
Answer by Theo(13342)

p = pamela's age.
j = jiri's age.

p = j - 10
p + j = 64
replace p with j - 10 to get:
j - 10 + j = 64
add 10 to both sides and combine like terms to get:
2j = 74
solve for j to get:
j = 37
that makes p = 27
27 + 37 = 64

solution is jiri's age is 37.